TFY4345 Classical Mechanics Autumn 2023

Course information

Homepage:

http://web.phys.ntnu.no/~stovneng/TFY4345_2023/tfy4345e.htm

Lecturer: Jon Andreas Støvneng Office: B4-141 Email: jon.stovneng@ntnu.no

Weekly schedule: See homepage.

Curriculum:

• Lectured material. See Lecture notes and Lecture plan.

• All assignments with solutions.

• Earlier exams are good training.

Textbook: Goldstein, Poole, Safko 3. edition (2002)

Examination arrangement: Written digital exam 27.11.23, 9-13. Counts 100 %. Examination aids C: Rottmann and allowed calculator.

Admission to exam:

• Approved assignments.

(Or: Approved assignments in TFY4345 from a previous year.)

Department of Physics, NTNU 11.08.2023

Jon Andreas Støvneng